OpenStudy (anonymous):

Solve for n. 7n = 35 A. 0.2 B. 5 C. 28 D. 42

OpenStudy (anonymous):

Its B. 35/7=5

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Which equation can be used to solve the problem? How many 12-packs of juice boxes contain a total of 84 juice boxes? A. 12 + b = 84 B. 12 ÷ b = 84 C. 12b = 84 D. b ÷ 12 = 84

OpenStudy (anonymous):

C right?

OpenStudy (anonymous):

B bruh 7 x 5 = 35